The charge flowing through a resistor varies with time as Q=2t-8t^2.The total heat produce in the resistor is?

Grade:12

5 Answers

Nicho priyatham
8 years ago
differenciate  and get i
now p=i2R
approve if useful in case of any doubt ask in ans box
wts the value of R
devajith
8 years ago
in the interva 0
devajith
8 years ago
sorry my comp hanged
in the interval 0
Resistance is R itelf
The answers are in fractions of R
devajith
8 years ago
(0,1/8)
 
Nicho priyatham
8 years ago
If Q = 2t - 8t2 then i = dq/dt
so i = 2-16t
so at t = 0 i = 2
since p = i2R  so p = 4R
plz approve if usefu n in case of any doubts put in ans box